# Quantum geometric tensor determines the i.i.d. conversion rate in the resource theory of asymmetry for any compact Lie group

### v3

@@ B.1 Proof of Lemma 3
$E_{p_{q},\rho}(A)=(1-q)\rho A+qA\rho=m_{f_{q}}(\mathcal{R}_{\rho},\mathcal{L}_{\rho})(A),$ (153)
where $m_{f}(x,y)\coloneqq yf(x/y)$ and $f_{q}$ is an operator monotone function defined in Eq. (24). Then, we have the following:
@@
B.2 Proof of the equivalence between two alternative definitions of asymptotic convertibility
@@ B.3 Proof of Lemma 4
Let $\sigma$ be an arbitrary state and $\phi$ be a pure state. Suppose that their infidelity by $\delta\coloneqq 1-\braket{\phi|\sigma|\phi}$ satisfies $\delta<1/2$. For the eigenvector $\Phi$ with the largest eigenvalue of $\sigma$, it holds
@@
$\braket{\mathrm{i}[\sigma,O],\mathrm{i}[\sigma,O]}_{f,\sigma}\geq\frac{1}{f\left(\frac{\delta}{1-\delta}\right)}(1-2\delta)^{2}\Braket{\Phi|O(I-\Phi)O^{\dagger}|\Phi}.$ (163)
@@
[…]
@@
As we have mentioned in Sec. IV.2, the variance per copy can generally change of the order of $O(N\epsilon_{N})$ for a state $\Phi_{N}$ that is $\epsilon_{N}$-close to $\phi^{\otimes N}$ in trace distance. Even when $\lim_{N\to\infty}\epsilon_{N}=0$, the variance per copy may change significantly in the limit of $N\to\infty$, and therefore, the variance has an asymptotic discontinuity. Nevertheless, Lemma 26 shows that when the state deviates slightly from $\phi^{\otimes N}$, the variance per copy can decrease at most $h_{1}(\epsilon)V(\phi,A)$, which is independent of $N$. The central limit theorem provides insight into this mechanism. The probability distribution of $\phi^{\otimes N}$ with respect to the observable $A_{N}$ is approximated by the normal distribution as $N\to\infty$ after properly redefining the random variable. When this probability distribution is modified due to the change in state, the variance cannot be reduced radically because the tail of the normal distribution is exponentially small.
@@
[…]
@@
$\Delta\mu_{N}$ $\coloneqq\braket{\phi^{\otimes N}|A_{N}|\phi^{\otimes N}}-\braket{\Phi_{N}|A_{N}|\Phi_{N}},$ (192)
$Q_{N}$ $\coloneqq(A_{N}-\braket{\Phi_{N}|A_{N}|\Phi_{N}}I)^{2}=(A_{N}-\braket{\phi^{\otimes N}|A_{N}|\phi^{\otimes N}}I+\Delta\mu_{N}I)^{2}.$ (193)
@@
$\braket{\phi^{\otimes N}|(A_{N}-\braket{\Phi_{N}}{A_{N}|\Phi_{N}}I)^{2}|\phi^{\otimes N}|=}\braket{\phi^{\otimes N}|(A_{N}-\braket{\phi^{\otimes N}}{A_{N}|\phi^{\otimes N}}I)^{2}|\phi^{\otimes N}|+}|\Delta\mu_{N}|^{2},$ (194)
@@
where $\mu\coloneqq\braket{\phi|A|\phi}$, $\sigma^{2}\coloneqq\braket{\phi|A^{2}|\phi}-\braket{\phi|A|\phi}^{2}$, and $x>0$ is a positive parameter that will be fixed later.
@@
From Lemmas 25 and 26, we obtain the following corollary:
@@
[…]
$\Braket{\Phi_{N}|O_{N}(I-\Phi_{N})O_{N}^{\dagger}|\Phi_{N}}\geq\braket{\phi^{\otimes N}|O_{N}(I-\phi^{\otimes N})O_{N}^{\dagger}|\phi^{\otimes N}}-Nh_{2}(\epsilon)$ (229)
@@
$\Braket{\Phi_{N}|O_{N}(I-\Phi_{N})O_{N}^{\dagger}|\Phi_{N}}\geq\braket{\phi^{\otimes N}|O_{N}(I-\phi^{\otimes N})O_{N}^{\dagger}|\phi^{\otimes N}}-Nh_{2}(\epsilon)$ (231)
@@
$\braket{\mathrm{i}[\sigma_{N},O],\mathrm{i}[\sigma_{N},O]}_{f,\sigma_{N}}\geq\frac{1}{f\left(\frac{\delta_{N}}{1-\delta_{N}}\right)}(1-2\delta_{N})^{2}\Braket{\Phi_{N}|O_{N}(I-\Phi)O_{N}^{\dagger}|\Phi_{N}},$ (234)
where $\delta_{N}\coloneqq 1-\braket{\phi^{\otimes N}|\sigma_{N}|\phi^{\otimes N}}$ denotes the infidelity. For a sufficiently small $\epsilon>0$, from Lemma 27, we further find
$\braket{\mathrm{i}[\sigma_{N},O],\mathrm{i}[\sigma_{N},O]}_{f,\sigma_{N}}\geq\frac{1}{f\left(\frac{\delta_{N}}{1-\delta_{N}}\right)}(1-2\delta_{N})^{2}\left(N\braket{\phi|O(I-\phi)O^{\dagger}|\phi}-Nh_{2}(\epsilon)\right)$ (235)
@@
$f(0)\braket{\mathrm{i}[\sigma_{N},O],\mathrm{i}[\sigma_{N},O]}_{f,\sigma_{N}}\geq\frac{f(0)}{f\left(\frac{\delta_{N}}{1-\delta_{N}}\right)}(1-2\delta_{N})^{2}N\braket{\phi|O(I-\phi)O^{\dagger}|\phi}-Nh(\epsilon),$ (236)
@@
$f(0)\braket{\mathrm{i}[\sigma_{N},O],\mathrm{i}[\sigma_{N},O]}_{f,\sigma_{N}}\geq N\braket{\phi|O(I-\phi)O^{\dagger}|\phi}-Nh(\epsilon)+N\braket{\phi|O(I-\phi)O^{\dagger}|\phi}\left(\frac{f(0)}{f\left(\frac{\delta_{N}}{1-\delta_{N}}\right)}(1-2\delta_{N})^{2}-1\right).$ (238)
@@
Appendix C Proofs of lemmas in the direct part
C.1 Proof of Lemma 7

### v4

@@ C.2.1 Proof of Lemma 3
$|\displaystyle E_{p_{q},\rho}(A)=(1-q)\rho A+qA\rho=\sum_{k,l}m_{f_{q}}(p_{k},p_{l})\ket{l}\braket{l|A|k}\bra{k}$ (140)
where $|\rho=\sum_{i=1}^{d}p_{i}\ket{i}\bra{i}$ is the eigenvalue decomposition of $\rho$, $m_{f}(x,y)\coloneqq yf(x/y)$ and $f_{q}$ is an operator monotone function defined in Eq. (25). Then, we have the following:
@@
C.2.2 Proof of the equivalence between two alternative definitions of asymptotic convertibility
@@ C.2.3 Proof of Lemma 17
[…]
@@
$\braket{\mathrm{i}[\sigma,O],\mathrm{i}[\sigma,O]}_{f,\sigma}\geq\frac{1}{f\left(\frac{\delta}{1-\delta}\right)}(1-2\delta)^{2}\mathrm{Tr}\left(\Phi O(I-\Phi)O^{\dagger}\right),$ (150)
where $|\Phi=\ket{\Phi}\bra{\Phi}$.
@@
Let $A$ be an arbitrary Hermitian operator on a finite-dimensional Hilbert space $\mathcal{H}$. We denote their i.i.d. extensions by $A_{N}$ defined in Eq. (168). Fix a pure state $\phi\in\mathcal{P}(\mathcal{H})$ and a sufficiently small parameter $\epsilon$. There exists a function $h_{1}(\epsilon)$ independent of $N$ satisfying $\lim_{\epsilon\to 0}h_{1}(\epsilon)=0$ such that for any sequence of pure states $\{\Phi_{N}\}_{N}$ satisfying $\limsup_{N\to\infty}T(\Phi_{N},\phi^{\otimes N})<\epsilon$, it holds
@@
As we have mentioned in Sec. IV.1, the variance per copy can generally change of the order of $O(N\epsilon_{N})$ for a state $\Phi_{N}$ that is $\epsilon_{N}$-close to $\phi^{\otimes N}$ in trace distance. Even when $\lim_{N\to\infty}\epsilon_{N}=0$, the variance per copy may change significantly in the limit of $N\to\infty$, and therefore, the variance has an asymptotic discontinuity. Nevertheless, Lemma 22 shows that when the state deviates slightly from $\phi^{\otimes N}$, the variance per copy can decrease at most $h_{1}(\epsilon)V(\phi,A)$, which is independent of $N$. The central limit theorem provides insight into this mechanism. The probability distribution of $\phi^{\otimes N}$ with respect to the observable $A_{N}$ is approximated by the normal distribution as $N\to\infty$ after properly redefining the random variable. When this probability distribution is modified due to the change in state, the variance cannot be reduced radically because the tail of the normal distribution is exponentially small.
@@
When all the eigenvalues of $A$ are degenerate, the statement trivially holds since $V(\Phi_{N},A_{N})=V(\phi^{\otimes N},A_{N})=0$. Below, we assume the difference between the largest and smallest eigenvalues of $A$ is non-vanishing, which we denote $\Delta a>0$. Similarly, if the variance $\sigma^{2}\coloneqq\mathrm{Tr}(\phi A^{2})-(\mathrm{Tr}(\phi A))^{2}$ vanishes for $\phi$, then the inequality is trivial since $V(\phi^{\otimes N},A_{N})=N\sigma^{2}=0$. Therefore, we also assume that $\sigma>0$.
@@
$\Delta\mu_{N}$ $\coloneqq\mathrm{Tr}\left(\phi^{\otimes N}A_{N}\right)-\mathrm{Tr}\left(\Phi_{N}A_{N}\right),$ (179)
$Q_{N}$ $\coloneqq\left(A_{N}-\mathrm{Tr}\left(\Phi_{N}A_{N}\right)I\right)^{2}=(A_{N}-\mathrm{Tr}\left(\phi^{\otimes N}A_{N}\right)I+\Delta\mu_{N}I)^{2}.$ (180)
@@
$\mathrm{Tr}\left(\phi^{\otimes N}(A_{N}-\mathrm{Tr}\left(\Phi_{N}A_{N}\right)I)^{2}\right)=\mathrm{Tr}\left(\phi^{\otimes N}(A_{N}-\mathrm{Tr}\left(\phi^{\otimes N}A_{N}\right)I)^{2}\right)+|\Delta\mu_{N}|^{2},$ (181)
@@
where $\mu\coloneqq\mathrm{Tr}(\phi A)$, $\sigma^{2}\coloneqq\mathrm{Tr}(\phi A^{2})-(\mathrm{Tr}(\phi A))^{2}$, and $x>0$ is a positive parameter that will be fixed later.
@@
From Lemmas 21 and 22, we obtain the following corollary:
@@
Let $O$ be an arbitrary linear operator on a finite-dimensional Hilbert space $\mathcal{H}$. We denote by $O_{N}$ its i.i.d. extension defined in Eq. (165). Fix a pure state $\phi\in\mathcal{P}(\mathcal{H})$ and a sufficiently small parameter $\epsilon$. Then there exists a function $h_{2}(\epsilon)$ independent of $N$ satisfying $\lim_{\epsilon\to 0}h_{2}(\epsilon)=0$ such that for any sequence of pure states $\{\Phi_{N}\}_{N}$ satisfying $\limsup_{N\to\infty}T(\Phi_{N},\phi^{\otimes N})<\epsilon$, it holds
$\mathrm{Tr}\left(\Phi_{N}O_{N}(I-\Phi_{N})O_{N}^{\dagger}\right)\geq\mathrm{Tr}\left(\phi^{\otimes N}O_{N}(I-\phi^{\otimes N})O_{N}^{\dagger}\right)-Nh_{2}(\epsilon)$ (216)
@@
$\mathrm{Tr}\left(\Phi_{N}O_{N}(I-\Phi_{N})O_{N}^{\dagger}\right)\geq\mathrm{Tr}\left(\phi^{\otimes N}O_{N}(I-\phi^{\otimes N})O_{N}^{\dagger}\right)-Nh_{2}(\epsilon)$ (218)
@@
$\braket{\mathrm{i}[\sigma_{N},O],\mathrm{i}[\sigma_{N},O]}_{f,\sigma_{N}}\geq\frac{1}{f\left(\frac{\delta_{N}}{1-\delta_{N}}\right)}(1-2\delta_{N})^{2}\mathrm{Tr}\left(\Phi_{N}O_{N}(I-\Phi_{N})O_{N}^{\dagger}\right),$ (221)
where $\delta_{N}\coloneqq 1-\mathrm{Tr}(\phi^{\otimes N}\sigma_{N})$ denotes the infidelity. For a sufficiently small $\epsilon>0$, from Lemma 23, we further find
$\braket{\mathrm{i}[\sigma_{N},O],\mathrm{i}[\sigma_{N},O]}_{f,\sigma_{N}}\geq\frac{1}{f\left(\frac{\delta_{N}}{1-\delta_{N}}\right)}(1-2\delta_{N})^{2}\left(N\mathrm{Tr}(\phi O(I-\phi)O^{\dagger})-Nh_{2}(\epsilon)\right)$ (222)
@@
$f(0)\braket{\mathrm{i}[\sigma_{N},O],\mathrm{i}[\sigma_{N},O]}_{f,\sigma_{N}}\geq\frac{f(0)}{f\left(\frac{\delta_{N}}{1-\delta_{N}}\right)}(1-2\delta_{N})^{2}N\mathrm{Tr}(\phi O(I-\phi)O^{\dagger})-Nh(\epsilon),$ (223)
@@
$f(0)\braket{\mathrm{i}[\sigma_{N},O],\mathrm{i}[\sigma_{N},O]}_{f,\sigma_{N}}\geq N\mathrm{Tr}(\phi O(I-\phi)O^{\dagger})-Nh(\epsilon)+N\mathrm{Tr}(\phi O(I-\phi)O^{\dagger})\left(\frac{f(0)}{f\left(\frac{\delta_{N}}{1-\delta_{N}}\right)}(1-2\delta_{N})^{2}-1\right).$ (225)
@@
Appendix D Proofs of the direct part
In this section, we present the proof of the direct part. In Sec. D.1, we provide a refined and detailed version of the arguments sketched in Sec. IV.2 of the main text. The proofs of the lemmas used in the direct part are given in Sec. D.2.
D.1 QGT determines asymptotic conversion rate
Proposition 24 (Direct part).
Let $U,U^{\prime}$ be (non-projective) unitary representations of a compact Lie group $G$ on finite-dimensional Hilbert spaces $\mathcal{H}$ and $\mathcal{H}^{\prime}$. For any pure states $\psi\in\mathcal{P}(\mathcal{H})$ and $\phi\in\mathcal{P}(\mathcal{H}^{\prime})$ such that $\mathrm{Sym}_{G}(\psi)\subset\mathrm{Sym}_{G}(\phi)$, it holds
Combining Proposition 24 with Eq. (38), our main result in Eq. (9) is proven for (non-projective) unitary representations, meaning that QGT determines the asymptotic conversion rate. As mentioned in Sec. III, this result can be extended to projective unitary representations by using the method in [64], which is detailed in Appendix B.3.
In the proof of Proposition 24, the results in QLAN on pure-state unitary models [72, 73] play a crucial role. This can be intuitively understood through Lemma 4, which provides an alternative yet equivalent characterization of the convertibility in RTA using quantum channels that are not assumed to be $G$-covariant.
D.1.1 Quantum local asymptotic normality and QGT
Here, let us briefly review studies in QLAN [70, 71, 72, 73], which investigate the asymptotic properties of statistical models by associating them with their limit models. This subsection discusses general statistical models, while the next subsection specializes them to the RTA setting.
We review particularly for QLAN on pure-state models [72, 73]. For a pure state $\psi$ and a set of Hermitian operators $\{X_{\mu}\}_{\mu=1}^{m}$, we define a pure-state statistical model $\mathcal{U}_{\theta}(\psi)$ with $\mathcal{U}_{\theta}(\cdot)=e^{\mathrm{i}\theta^{\mu}X_{\mu}}(\cdot)e^{-\mathrm{i}\theta^{\nu}X_{\nu}}$ for $\theta\in\mathbb{R}^{m}$. To study its asymptotic property, we consider its i.i.d. copies of the state $(\mathcal{U}_{u/\sqrt{N}}(\psi))^{\otimes N}$, where we introduced a local parameter $u\coloneqq\sqrt{N}\theta$. In the limit of $N\to\infty$, it is shown [72, 73] that this model is interconvertible to the so-called Gaussian shift model with asymptotically vanishing error.
[…]
$=0$ (228)
for $\epsilon\in(0,1/9)$.
[…]
Importantly, the $(\mu,\nu)$ element of $C^{\dagger}C$ is given by
$\left(C^{\dagger}C\right)_{\mu\nu}$ $=\sum_{k=1}^{d-1}\braket{\psi|X_{\mu}|k}\braket{k|X_{\nu}|\psi}$
$=\braket{\psi|X_{\mu}(I-\psi)X_{\nu}|\psi},$ (229)
meaning that $C^{\dagger}C=\mathcal{Q}^{\psi}$, where $\mathcal{Q}^{\psi}$ is the QGT for $\mathcal{U}_{\theta}(\psi)$. Therefore, if two pure-state statistical models have the same QGT, they are asymptotically interconvertible to each other since both of them can be reversibly converted to the same Gaussian shift model, shown in Eq. (228).
As a generalization of the above observation, we prove the following:
Lemma 25.
For sets of Hermitian operators $\{X_{\mu}\}_{\mu=1}^{m}$ and $\{X_{\mu}^{\prime}\}_{\mu=1}^{m}$, we define $\mathcal{U}_{\theta}(\cdot)\coloneqq e^{\mathrm{i}\theta^{\mu}X_{\mu}}(\cdot)e^{-\mathrm{i}\theta^{\nu}X_{\nu}}$ and $\mathcal{U}^{\prime}_{\theta}(\cdot)\coloneqq e^{\mathrm{i}\theta^{\mu}X_{\mu}^{\prime}}(\cdot)e^{-\mathrm{i}\theta^{\nu}X_{\nu}^{\prime}}$ for $\theta\in\mathbb{R}^{m}$. For given two pure states $\psi$ and $\phi$, we define pure-state statistical models by $\mathcal{U}_{\theta}(\psi)$ and $\mathcal{U}^{\prime}_{\theta}(\phi)$ and denote their QGTs by $\mathcal{Q}^{\psi}$ and $\mathcal{Q}^{\phi}$, respectively. If $r>0$ satisfies $\mathcal{Q}^{\psi}\geq r\mathcal{Q}^{\phi}$, then there exists a sequence of quantum channels $\{\mathcal{E}_{N}\}_{N}$ such that
$\lim_{N\to\infty}\sup_{\|u\|<N^{\epsilon}}T\left(\mathcal{E}_{N}\left(\mathcal{U}_{\frac{u}{\sqrt{N}}}(\psi)^{\otimes N}\right),\mathcal{U}^{\prime}_{\frac{u}{\sqrt{N}}}(\phi)^{\otimes\lfloor rN\rfloor}\right)=0$ (230)
for $\epsilon\in(0,1/9)$.
In the proof of Lemma 25, the argument is divided into two parts: one focuses on the case where the QGT is conserved, while the other addresses the case where the QGT is reduced. We first prove that the growth rate of i.i.d. copies can be set to one by adjusting the scaling of the parameters. Concretely, $\mathcal{U}^{\prime}_{\frac{u}{\sqrt{N}}}(\phi)^{\otimes\lfloor rN\rfloor}$ is shown to be asymptotically interconvertible to $\mathcal{U}^{\prime}_{\frac{\sqrt{r}u}{\sqrt{N}}}(\phi)^{\otimes N}$. This asymptotic interconversion is achieved by combining the reversible conversion channels in QLAN, where the QGT is conserved. Then, we also prove that $\mathcal{U}_{\frac{u}{\sqrt{N}}}(\psi)^{\otimes N}$ can be asymptotically converted into $\mathcal{U}^{\prime}_{\frac{\sqrt{r}u}{\sqrt{N}}}(\phi)^{\otimes N}$ if the QGT with respect to $u$ decreases in the sense of matrix inequality. The details of these lemmas and their proofs are provided in Appendix D.2.2.
D.1.2 Proof of achievability of the conversion rate
The process of transforming $\mathcal{U}_{g}(\psi)^{\otimes N}$ to $\mathcal{U}_{g}^{\prime}(\phi)^{\otimes\lfloor rN\rfloor}$ with asymptotically vanishing error consists of two steps, which we call the estimation and conversion steps. In the estimation step, we obtain a rough estimate $\hat{g}$ of the true value $g$ by consuming a sublinear number of copies of $\mathcal{U}_{g}(\psi)$ in $N$. In the conversion step, based on the estimate $\hat{g}$, we convert the remaining copies of $\mathcal{U}_{g}(\psi)$ to i.i.d. copies of $\mathcal{U}_{g}^{\prime}(\phi)$ with quantum channels constructed by using Lemma 25.
For a detailed explanation, let us first introduce notations. Let $U$ and $U^{\prime}$ be (non-projective) unitary representations of a Lie group $G$ on the input and output systems. Elements in the neighborhood of the identity $e\in G$ can be parametrized as $g(\lambda)=e^{\mathrm{i}\lambda^{\mu}A_{\mu}}$ by using a basis $\{A_{\mu}\}_{\mu=1}^{\dim G}$ of the Lie algebra $\mathfrak{g}$. Define operators $X_{\mu}\coloneqq-\mathrm{i}\frac{\partial}{\partial\lambda^{\mu}}U(g(\lambda))|_{\lambda=0}$ and $X^{\prime}_{\mu}\coloneqq-\mathrm{i}\frac{\partial}{\partial\lambda^{\mu}}U^{\prime}(g(\lambda))|_{\lambda=0}$, which correspond to the images of $A_{\mu}$ by the Lie algebra representations induced from $U$ and $U^{\prime}$. At least locally, these operators satisfy $U(g(\lambda))=e^{\mathrm{i}\lambda^{\mu}X_{\mu}}$ and $U^{\prime}(g(\lambda))=e^{\mathrm{i}\lambda^{\mu}X_{\mu}^{\prime}}$. In this case, by slightly abusing our notation of $\mathcal{U}_{g}$ and $\mathcal{U}^{\prime}_{g}$, we also denote $\mathcal{U}_{\theta}(\cdot)\coloneqq e^{\mathrm{i}\theta^{\mu}X_{\mu}}(\cdot)e^{-\mathrm{i}\theta^{\nu}X_{\nu}}$ and $\mathcal{U}_{\theta}^{\prime}(\cdot)\coloneqq e^{\mathrm{i}\theta^{\mu}X_{\mu}^{\prime}}(\cdot)e^{-\mathrm{i}\theta^{\nu}X_{\nu}^{\prime}}$ for $\theta\in\mathbb{R}^{\dim G}$, implicitly indicating $\|\theta\|$ is sufficiently small.
$G_{\mathrm{succ.}}^{(g,\delta)}\coloneqq\{\hat{g}\in G\mid\exists\theta,\,\mathcal{U}_{\theta}\circ\mathcal{U}_{\hat{g}}(\psi)=\mathcal{U}_{g}(\psi),\,\|\theta\|<\delta\}.$ (231)
We quantify the success probability in the worst case by
[…]
where $\mu_{G}$ is the Haar measure on the Lie group $G$, and $\chi_{A}$ denotes the indicator function of a set $A$. Note that the above success probability is well-defined since $G_{\mathrm{succ.}}^{(g,\delta)}$ is a measurable set, as shown in Appendix D.2.3.
Lemma 26.
$\lim_{N\to\infty}p^{\mathrm{succ.}}(N^{-1/2+\epsilon})=1.$ (233)
We remark that $\|\theta\|\sim N^{-1/2}$ for the error is the standard scaling in parameter estimation for i.i.d. models [135], including QLAN studies [70, 71, 72, 73]. The proof is detailed in Appendix D.2.4.
Lemma 27.
Let $U,U^{\prime}$ be (non-projective) unitary representations of a compact Lie group $G$ on finite-dimensional Hilbert spaces $\mathcal{H}$ and $\mathcal{H}^{\prime}$. Let $\psi\in\mathcal{P}(\mathcal{H})$ and $\phi\in\mathcal{P}(\mathcal{H}^{\prime})$ be pure states. If $r>0$ satisfies $\mathcal{Q}^{\mathcal{U}_{g}(\psi)}\geq r\mathcal{Q}^{\mathcal{U}_{g}^{\prime}(\phi)}$ for all $g\in G$, then, there exists a sequence of quantum channels $\{\mathcal{E}^{(g)}_{N}\}_{N}$ such that the conversion error
$\delta_{N}(g,u)\coloneqq T\left(\mathcal{E}^{(g)}_{N}\left(\mathcal{U}_{\frac{u}{\sqrt{N}}}\left(\mathcal{U}_{g}(\psi)\right)^{\otimes N}\right),\mathcal{U}^{\prime}_{\frac{u}{\sqrt{N}}}\left(\mathcal{U}_{g}^{\prime}(\phi)\right)^{\otimes\lfloor rN\rfloor}\right)$ (234)
satisfies $\lim_{N\to\infty}\sup_{g\in G}\sup_{\|u\|<N^{\epsilon}}\delta_{N}(g,u)=0$ for $\epsilon\in(0,1/9)$.
This lemma is proven from Lemma 25 by using properties of a compact Lie group. The proof is detailed in Appendix D.2.5.
Proposition 24 is proven by combining Lemmas 26 and 27.
Proof of Proposition 24.
$\mathcal{E}_{N}(\mathcal{U}_{g}(\psi)^{\otimes N})\coloneqq\int\mathrm{d}\mu_{G}(\hat{g})\,\,p(\hat{g}|\mathcal{U}_{g}(\psi)^{\otimes\lceil N^{1-\epsilon}\rceil})\mathcal{E}_{N^{\prime}}^{(\hat{g})}\left(\mathcal{U}_{g}(\psi)^{\otimes N^{\prime}}\right).$ (235)
From the assumption $\mathrm{Sym}_{G}(\psi)\subset\mathrm{Sym}_{G}(\phi)$, if parameters $\theta\in\mathbb{R}^{\dim G}$ satisfy $\mathcal{U}_{\theta}\circ\mathcal{U}_{\hat{g}}(\psi)=\mathcal{U}_{g}(\psi)$ for some $g,\hat{g}\in G$, then it also holds $\mathcal{U}^{\prime}_{\theta}\circ\mathcal{U}^{\prime}_{\hat{g}}(\phi)=\mathcal{U}_{g}^{\prime}(\phi)$. Therefore, for $\hat{g}\in G_{\mathrm{succ.}}^{(g,\delta)}$ with $\delta>0$, we have
$\mathrm{Fid}\left(\mathcal{E}_{N^{\prime}}^{(\hat{g})}\left(\mathcal{U}_{g}(\psi)^{\otimes N^{\prime}}\right),\mathcal{U}_{g}^{\prime}(\phi)^{\otimes\lfloor rN^{\prime}\rfloor}\right)=\mathrm{Fid}\left(\mathcal{E}_{N^{\prime}}^{(\hat{g})}\left(\mathcal{U}_{\theta}(\mathcal{U}_{\hat{g}}(\psi))^{\otimes N^{\prime}}\right),\mathcal{U}_{\theta}^{\prime}(\mathcal{U}_{\hat{g}}^{\prime}(\phi))^{\otimes\lfloor rN^{\prime}\rfloor}\right),$ (236)
where $\mathrm{Fid}$ denotes the fidelity defined by $\mathrm{Fid}(\rho,\sigma)\coloneqq\left(\mathrm{Tr}\left(\sqrt{\sqrt{\rho}\sigma\sqrt{\rho}}\right)\right)^{2}$ for states $\rho$ and $\sigma$. Since $N^{-1/2+\epsilon}<(N^{\prime})^{-1/2+\epsilon}$, we have
$f_{N}$ $\coloneqq\inf_{g\in G}\inf_{\hat{g}\in G_{\mathrm{succ.}}^{(g,N^{-1/2+\epsilon})}}\mathrm{Fid}\left(\mathcal{E}_{N^{\prime}}^{(\hat{g})}\left(\mathcal{U}_{g}(\psi)^{\otimes N^{\prime}}\right),\mathcal{U}_{g}^{\prime}(\phi)^{\otimes\lfloor rN^{\prime}\rfloor}\right)$ (237)
$=\inf_{\hat{g}\in G}\inf_{\|u\|<(N^{\prime})^{\epsilon}}\mathrm{Fid}\left(\mathcal{E}_{N^{\prime}}^{(\hat{g})}\left(\mathcal{U}_{\frac{u}{\sqrt{N^{\prime}}}}(\mathcal{U}_{\hat{g}}(\psi))^{\otimes N^{\prime}}\right),\mathcal{U}_{\frac{u}{\sqrt{N^{\prime}}}}^{\prime}(\mathcal{U}^{\prime}_{\hat{g}}(\phi))^{\otimes\lfloor rN^{\prime}\rfloor}\right)$
$\geq\left(1-\sup_{\hat{g}\in G}\sup_{\|u\|<(N^{\prime})^{\epsilon}}\delta_{N^{\prime}}(\hat{g},u)\right)^{2}\xrightarrow{N\to\infty}1,$ (238)
where in the last line, we used the Fuchs–van de Graaf inequalities to relate the fidelity to the trance distance $\delta_{N}(g,u)$ defined in Eq. (234) in Lemma 27.
From Eqs. (233), (235), and (238), we get
$\inf_{g\in G}\mathrm{Fid}\left(\mathcal{E}_{N}(\mathcal{U}_{g}(\psi)^{\otimes N}),\mathcal{U}_{g}^{\prime}(\phi)^{\otimes\lfloor rN^{\prime}\rfloor}\right)$ $=\inf_{g\in G}\int_{G}\mathrm{d}\mu_{G}(\hat{g})\,\,p(\hat{g}|\mathcal{U}_{g}(\psi)^{\otimes\lceil N^{1-\epsilon}\rceil})$
[…]
$\geq p^{\mathrm{succ.}}(N^{-1/2+\epsilon})\times f_{N}$ (239)
$\xrightarrow{N\to\infty}1,$ (240)
implying that
From Lemma 4, this is equivalent to $\{\psi^{\otimes N}\}_{N}\xrightarrow{G\text{-cov.}}\{\phi^{\otimes\lfloor(r-\delta)N\rfloor}\}_{N}$, which completes the proof of Proposition 24. ∎
The above proof of the direct part provides a clear understanding of the reason why the asymptotic conversion rate diverges for finite groups. Since there is only a finite number of different elements in $\{\mathcal{U}_{g}(\psi)\mid g\in G\}$, it is possible to identify the state $\mathcal{U}_{g}(\psi)$ with an exponentially small failure probability using a state tomography protocol on $\mathcal{U}_{g}(\psi)^{\otimes N}$. Let $\hat{g}\in G$ denote the estimate of the true value $g$ obtained by this tomography protocol. Depending on the estimate $\hat{g}$, we can prepare an arbitrary number of copies of $\mathcal{U}^{\prime}_{\hat{g}}(\phi)$. The condition $\mathrm{Sym}_{G}(\psi)\subset\mathrm{Sym}_{G}(\phi)$ implies that if $\mathcal{U}_{\hat{g}}(\psi)=\mathcal{U}_{g}(\psi)$ is satisfied, then $\mathcal{U}^{\prime}_{\hat{g}}(\phi)=\mathcal{U}^{\prime}_{g}(\phi)$ also holds. Therefore, the asymptotic conversion rate diverges since it is possible to create an arbitrary number of copies of $\mathcal{U}^{\prime}_{g}(\phi)$ from $\mathcal{U}_{g}(\psi)^{\otimes N}$ with asymptotically vanishing failure probability. We emphasize that the exact asymptotic conversion rate in [64] is not covered by this analysis because there always remains an error in the tomography process.
D.2 Proof of Lemmas in the direct part
D.2.1 Proof of Lemma 4
For $\epsilon=0$, i.e., for the conversion without error, Lemma 4 is shown as Lemma 7 in [14]. Although its generalization to the case of $\epsilon>0$ is straightforward, we here provide the proof for completeness.